# LLM as BT-Planner: Leveraging LLMs for Behavior Tree Generation in Robot Task Planning

### v1

@@ Abstract
Robotic assembly tasks are open challenges due to the long task horizon and complex part relations. Behavior trees (BTs) are increasingly used in robot task planning for their modularity and flexibility, but manually designing them can be effort-intensive. Large language models (LLMs) have recently been applied in robotic task planning for generating action sequences, but their ability to generate BTs has not been fully investigated. To this end, We propose LLM as BT-planner, a novel framework to leverage LLMs for BT generation in robotic assembly task planning and execution. Four in-context learning methods are introduced to utilize the natural language processing and inference capabilities of LLMs to produce task plans in BT format, reducing manual effort and ensuring robustness and comprehensibility. We also evaluate the performance of fine-tuned, fewer-parameter LLMs on the same tasks. Experiments in simulated and real-world settings show that our framework enhances LLMs’ performance in BT generation, improving success rates in BT generation through in-context learning and supervised fine-tuning.
@@ I INTRODUCTION
Sequential manipulation planning has been a critical imperative to achieve a higher level of autonomy in robotics. Classical approaches to address task planning problems, such as Planning Domain Definition Language (PDDL) [1], are based on symbolic formalisms to search for transition paths in the state space to reach task goals. In practice, such task plans are often programmed as Finite State Machines (FSMs), which incorporate expert knowledge specifying control and execution details. Due to its limitation in scalability [2], behavior trees (BTs), which represent policies in a state-implicit, hierarchical tree structure, have gained increasing popularity for complex task planning. Its advantages of modularity, reusability, and reactivity make it a more desired formalism for long-horizon manipulation tasks.
Despite being more efficient to program, maintain, and modify than FSMs [3], manually programming BTs still requires significant effort. To automate the generation of BT-based task plans in the context of sequential robot manipulation, progress has been made by using (i) symbolic planning [4], (ii) learning from demonstration [5], and (iii) reinforcement learning [6]. However, these methods still largely rely on manual design for the basic structure or available subtrees of BTs. Methods for further automating generation still need to be further explored. .
@@
To this end, we propose LLM as BT-planner, an LLM-based BT generation framework to leverage the strengths of both for sequential manipulation planning, in which different approches are applied to enhance BT generation performance. To enable human-robot collaborative task planning and enhance intuitive robot programming by nonexperts, the framework takes human instructions to initiate subgoal sequence generation and human feedback to refine BT generation in runtime. Moreover, the improvement of LLMs’ performance in generating BTs after fine-tuning is investigated based on two different task types. The presented framework is tested on a real robotic assembly use case, which uses a gear set model from the Siemens Robot Assembly Challenge. We use a single manipulator with a tool-changing mechanism, a common practice in flexible manufacturing, to facilitate robust grasping of a large variety of objects.
@@
    We established a novel framework to generate fully executable BTs using LLMs in complex robotic assembly tasks with a set of performance metrics for evaluation. To the best of our knowledge, this is the first unified framework to leverage LLMs through the whole BT generation pipeline: (i) task decomposition from language instructions to sequence of subgoals, (ii) converting subgoals to parameterized task plans, (iii) generating BTs given task plans.
@@
    The performance of the proposed methods are evaluated in a systematic way. Our experiment results demonstrate that the human-in-the-loop approach outperforms other in-context learning methods. While fine-tuning improves the executability of generated BTs, reflecting better in-context understanding and structural generation, it has minimal effect on boosting the success rate of BT generation for smaller (compared to GPT-4) LLMs.
@@ V-A Experiment Setup
The experiment setup of physical robot validation is shown in Fig. 5. Operation space control is applied to a single 7-DoF Franka Emika Panda robot arm via Franka Control Interface (FCI) [40]. The computed torques come from a Cartesian adaptive force-impedance controller. To execute the actions in BTs, the action context is parsed to another custom control software, i.e. the skill base, to map the action to a pre-defined skill. PyTrees and WebSocket are used to implement asynchronous BTs. The world model integrates a self-implemented relation graph with Neo4j for world state management and visualization.
@@
The details of the use case are shown in Fig. 5. In the experiment, the robot is provided with tool cubes classified as parallelgripper, clampgripper, outwardgripper, and inwardgripper according to their self-designed fingertips. The generation methods are evaluated using the Siemens Robotic Assembly Challenge gearset use case, involving a gear base, three shafts, and three gears. The task is to assemble the shafts into the gearbase and insert the gears into their corresponding shafts to mesh. Here, the second shaft fixed on the gearbase is modified to be screwable to diverse the necessary robot skill types. Different LLMs are employed for BT generation across all four proposed approaches. For validation, the human-in-the-loop generation method is adopted for its superior performance, and the experiment convenience its human feedback brings.
@@ V-B Fine-tuning Process
Two open-source LLMs, Mistral-7B and Llama2-13B-chat, are selected and fine-tuned for the two task types. Besides, a GPT-3.5 model is also fine-tuned to show the performance of LLMs with larger parameter amounts. For the unit-tree generation task, training data are collected from the roll-out records of the recursive generation method with Siemens’ gearset assembly task as the use case. For the one-step generation task, the data collected from the one-step generation method are adopted as training data. In the training phase, both Mistral-7B and Llama2-13B-chat are trained using the Llama-factory framework [42]. For unit-tree generation tasks, Mistral-7B and Llama2-13B-chat are trained for 10 epochs with a learning rate of $1\times 10^{-4}$, while for One-step generation tasks, they are trained for 15 epochs with a learning rate of $5\times 10^{-5}$. GPT-3.5 is fine-tuned with a learning rate multiplier of 0.05 for 3 epochs for unit-tree generation tasks, while for one-step generation tasks, the same learning rate multiplier is applied for 10 epochs. After training, the LLMs’ performance on both task types is validated with the data collected from the chair and lamp use cases of the furniture assembly benchmark [41].
@@ V-D In-context Learning Results
Specific and directed user feedback significantly optimizes LLMs’ performance. The human-in-the-loop method demonstrates a significant improvement in logical coherence (16/17 compared to 12/17 of the one-step method) and excitability (17/17, better than 13/17 of the recursive method) because of the incorporation of precise user feedback, which also leads to larger average generation duration (85.02 seconds) and token consumption (7483.34 tokens). Notably, this method also allows the introduction of new natural language-represented knowledge for BT generation. The LLM leverages its powerful natural language processing capabilities to prioritize new input, such as changes in tool-object compatibility, and generates BTs based on this information rather than relying on outdated knowledge. This allows for efficient replanning without altering the knowledge base.
Algorithm guiding enhances BT generation performance with unstable structured output performance and more resource consumption. The recursive method, while ensuring high logical coherence (17/17, the highest among all methods), incurs the largest resource consumption (231.04 seconds for generating duration and 50229.96 for token consumption), reflecting its thoroughness in distributing the generation task across multiple recursive LLM invokes. These results highlight this method’s trade-offs between generation time, complexity, and accuracy. Furthermore, it also shows the worst excitability result (13/17). This is caused by the recursive invocations in the generation process, which magnifies the instability of LLMs’ generation capability.
@@ VI CONCLUSIONS
This work introduces LLM as BT-planner, an innovative framework that leverages LLMs for BT generation in robotic task planning. Various BT generation methods are explored based on different LLMs using in-context learning and fine-tuning techniques. Experimental evaluations show that our framework effectively generates BTs for robotic assembly tasks, offering a promising solution for complex task planning scenarios. Among the four BT generation methods, the human-in-the-loop approach outperforms the others. Fine-tuning improves LLMs’ performance in context understanding and structural generation but has an insignificant impact on few-parameter LLMs inference capability. This may be due to insufficient training data and limitations in model parameters, which need to be further investigated in future work.

### v2

# LLM-as-BT-Planner: Leveraging LLMs for Behavior Tree Generation in Robot Task Planning

@@ Abstract
Robotic assembly tasks remain an open challenge due to their long horizon nature and complex part relations. Behavior trees (BTs) are increasingly used in robot task planning for their modularity and flexibility, but creating them manually can be effort-intensive. Large language models (LLMs) have recently been applied to robotic task planning for generating action sequences, yet their ability to generate BTs has not been fully investigated. To this end, we propose LLM-as-BT-Planner, a novel framework that leverages LLMs for BT generation in robotic assembly task planning. Four in-context learning methods are introduced to utilize the natural language processing and inference capabilities of LLMs for producing task plans in BT format, reducing manual effort while ensuring robustness and comprehensibility. Additionally, we evaluate the performance of fine-tuned smaller LLMs on the same tasks. Experiments in both simulated and real-world settings demonstrate that our framework enhances LLMs’ ability to generate BTs, improving success rate through in-context learning and supervised fine-tuning.
@@ I INTRODUCTION
Sequential manipulation planning has been a critical imperative to achieve a higher level of autonomy in robotics. Classical approaches to address task planning problems, such as Planning Domain Definition Language (PDDL) [1], are based on symbolic formalisms to search for transition paths in the state space to reach task goals. In practice, such task plans are often programmed as Finite State Machines, which incorporate expert knowledge specifying control and execution details. Due to its limitation in scalability [2], behavior trees (BTs), which represent policies in a state-implicit, hierarchical tree structure, have gained increasing popularity for complex task planning. Its advantages of modularity, reusability, and reactivity make it a more desired formalism for long-horizon manipulation tasks.
Despite being more efficient to program, maintain, and modify than Finite State Machines [3], manually programming BTs still requires significant effort. To automate the generation of BT-based task plans in the context of sequential robot manipulation, progress has been made by using (i) symbolic planning [4], (ii) learning from demonstration [5], and (iii) reinforcement learning [6]. However, these methods still largely rely on manual design for the basic structure or available subtrees of BTs. Methods for further automating generation still need to be further explored. .
@@
To this end, we propose LLM-as-BT-planner, an LLM-based BT generation framework to leverage the strengths of both for sequential manipulation planning, in which different approaches are applied to enhance BT generation performance. To enable human-robot collaborative task planning and enhance intuitive robot programming by nonexperts, the framework takes human instructions to initiate subgoal sequence generation and human feedback to refine BT generation in runtime. Moreover, the improvement of LLMs’ performance in generating BTs after fine-tuning is investigated based on two different task types. The presented framework is tested on a real robotic assembly use case, which uses a gearset model from the Siemens Robot Assembly Challenge. We use a single manipulator with a tool-changing mechanism, a common practice in flexible manufacturing, to facilitate robust grasping of a large variety of objects.
@@
    We established a novel framework to generate fully executable BTs using LLMs in complex robotic assembly tasks with a set of performance metrics for evaluation. The BT generation pipeline includes (i) task decomposition from language instructions to sequence of subgoals, (ii) converting subgoals to parameterized task plans, and (iii) generating BTs given task plans.
@@
    The performance of the proposed methods is evaluated in a systematic way. Our experiment results demonstrate that the human-in-the-loop approach outperforms other in-context learning methods. While fine-tuning improves the executability of generated BTs, reflecting better in-context understanding and structural generation, it has minimal effect on boosting the success rate of BT generation for smaller (compared to GPT-4) LLMs.
@@ V-A Experiment Setup
The experiment setup of physical robot validation is shown in Fig. 5. Operation space control is applied to a single 7-DoF Franka Emika Panda robot arm via Franka Control Interface (FCI) [40]. The computed torques come from a Cartesian adaptive force-impedance controller. To execute the actions in BTs, the action context is parsed to another custom control software, i.e., the skill base, to map the action to a pre-defined skill. PyTrees and WebSocket are used to implement asynchronous BTs. The world model integrates a self-implemented relation graph with Neo4j for world state management and visualization.
@@
The details of the use case are shown in Fig. 5. In the experiment, the robot is provided with tool cubes classified as parallelgripper, clampgripper, outwardgripper, and inwardgripper according to their self-designed fingertips. The generation methods are evaluated using the Siemens Robotic Assembly Challenge gearset use case, involving a gear base, three shafts, and three gears. The task is to assemble the shafts into the gearbase and insert the gears into their corresponding shafts to mesh. Here, the second shaft fixed on the gearbase is modified to be screwable to diverse the necessary robot skill types. Different LLMs are employed for BT generation across all four proposed approaches. For validation, the human-in-the-loop generation method is adopted for its superior performance and the experiment convenience its human feedback brings.
@@ V-B Fine-tuning Process
Two open-source LLMs, Mistral-7B and Llama2-13B-chat, are selected and fine-tuned for the two task types. A GPT-3.5 model is also fine-tuned to show the performance of LLMs with larger parameter amounts. For the unit-tree generation task, training data are collected from the roll-out records of the recursive generation method with Siemens’ gearset assembly task as the use case. For the one-step generation task, the data collected from the one-step generation method are adopted as training data. In the training phase, both Mistral-7B and Llama2-13B-chat are trained using the Llama-factory framework [42]. For unit-tree generation tasks, Mistral-7B and Llama2-13B-chat are trained for 10 epochs with a learning rate of $1\times 10^{-4}$, while for One-step generation tasks, they are trained for 15 epochs with a learning rate of $5\times 10^{-5}$. GPT-3.5 is fine-tuned with a learning rate multiplier of 0.05 for 3 epochs for unit-tree generation tasks, while for one-step generation tasks, the same learning rate multiplier is applied for 10 epochs. After training, the LLMs’ performance on both task types is validated with the data collected from the chair and lamp use cases of the furniture assembly benchmark [41].
@@ V-D In-context Learning Results
Specific and directed user feedback significantly optimizes LLMs’ performance. The human-in-the-loop method demonstrates a significant improvement in logical coherence (16/17 compared to 12/17 of the one-step method) and executability (17/17, better than 13/17 of the recursive method) because of the incorporation of precise user feedback, which also leads to larger average generation duration (85.02 seconds) and token consumption (7483.34 tokens). Notably, this method also allows the introduction of new natural language-represented knowledge for BT generation. The LLM leverages its powerful natural language processing capabilities to prioritize new input, such as changes in tool-object compatibility, and generates BTs based on this information rather than relying on outdated knowledge. This allows for efficient replanning without altering the knowledge base.
Algorithm guiding enhances BT generation performance with unstable structured output performance and more resource consumption. The recursive method, while ensuring high logical coherence (17/17, the highest among all methods), incurs the largest resource consumption (231.04 seconds for generating duration and 50229.96 for token consumption), reflecting its thoroughness in distributing the generation task across multiple recursive LLM invokes. These results highlight this method’s trade-offs between generation time, complexity, and accuracy. Furthermore, it also shows the worst executability result (13/17). This is caused by the recursive invocations in the generation process, which magnifies the instability of LLMs’ generation capability.
@@ VI CONCLUSIONS
This work introduces LLM-as-BT-planner, an innovative framework that leverages LLMs for BT generation in robotic task planning. Various BT generation methods are explored based on different LLMs using in-context learning and fine-tuning techniques. Experimental evaluations show that our framework effectively generates BTs for robotic assembly tasks, offering a promising solution for complex task planning scenarios. Among the four BT generation methods, the human-in-the-loop approach outperforms the others. Fine-tuning improves LLMs’ performance in context understanding and structural generation but has an insignificant impact on few-parameter LLMs inference capability. This may be due to insufficient training data and limitations in model parameters, which need to be further investigated in future work.
ACKNOWLEDGMENT
The authors acknowledge the financial support by the Bavarian State Ministry for Economic Affairs, Regional Development and Energy (StMWi) for the Lighthouse Initiative KI.FABRIK (Phase 1: Infrastructure as well as the research and development program under, grant no. DIK0249). The authors acknowledge the financial support by the Federal Ministry of Education and Research of Germany (BMBF) in the programme of ”Souverän. Digital. Vernetzt.” Joint project 6G-life, project identification number 16KISK002.